OpenStudy (anonymous):

The sum of the three numbers is -4. The second number decreased by the third is equal to the first. The sum of the first and second numbers is -5. Find the numbers. a. 2, 5, 6. b. -2, 2, 3. c. 6, 7, -9. d. -3, -2, 1.

OpenStudy (anonymous):

Do you have an idea of what it might be? I don't want to give you the answer, but I'll definitely help you find it.

OpenStudy (aum):

The quickest way to solve this problem is to go through each answer choice and eliminate the wrong ones: "The sum of the first and second numbers is -5" Go through each answer choice and see which choice has the first and second number add up to -5. Just to be safe make sure the other conditions given in the problem are also satisfied: "The sum of the three numbers is -4" and "The second number decreased by the third is equal to the first."

OpenStudy (anonymous):

Without possible answers - construct system of linear equations and solve it: a+b+c=-4 b-c=a a+b=-5 http://www.hackmath.net/en/calculator/solving-system-of-linear-equations?input=a%2Bb%2Bc%3D-4%0D%0Ab-c%3Da%0D%0Aa%2Bb%3D-5&submit=Calculate Result: a = -3 b = -2 c = 1 do D is good